OpenStudy (anonymous):

Complete the square to form a perfect square trinomial. x² + 6x + __

OpenStudy (saifoo.khan):

How?

OpenStudy (karatechopper):

o wait oops

OpenStudy (saifoo.khan):

lol

OpenStudy (karatechopper):

wait no i think im rite!

OpenStudy (karatechopper):

ugh im confuzzled

OpenStudy (anonymous):

i have 9 as an option

OpenStudy (anonymous):

ax^2+bx+c for it to be a perfect square trinomial, you want to factor out the a. then take half of the coefficient "b" and square it a=1, so no need to factor it out x^2+6x+c c=(6/2)^2=3^2=9

OpenStudy (karatechopper):

ha! see im rite! so it is 9!

hero (hero):

lol

OpenStudy (saifoo.khan):

I just asked for method Kc. I never said u were wrong.

OpenStudy (karatechopper):

argghhhhhhhhhh bro u confuze me 2 much

OpenStudy (saifoo.khan):

hehe/ ;)

hero (hero):

KC, if you really understand it, no one can confuse you

OpenStudy (karatechopper):

u see im multitaskin here!!!